If the diameter of the circle above is 6 cm, what is the area of the circle?
Answer by only4christ(12)

Area = pi*r^2
Diameter = 2*r so r = d/2
Given: Diameter=6; dividing by 2 we get r; so r = 3
Area = 3.14 * (3)^2
A = 3.14 * 9=28.26 sq.cm.